# The Hofstadter consecutive-sum sequence omits infinitely many positive integers

### 10 Mar 2026 v1

@@ 1. Introduction
A self-generating greedy process, often discussed under the name Hofstadter sequences (see, for instance, [9] or [6, E31]), starts from a short initial list and then repeatedly adjoins sums of consecutive earlier terms. In its classical form, the seed is $(1,2)$, and one generates a strictly increasing sequence $(a_{n})_{n\geq 1}$ by setting $a_{1}=1$, $a_{2}=2$, and, for $k\geq 3$, taking $a_{k}$ to be the least integer greater than $a_{k-1}$ that can be written as a sum of at least two consecutive earlier terms:
@@
This is A005243 in the OEIS [8]. The asymptotic behavior of $(a_{n})_{n\geq 1}$ was asked about by Hofstadter and has been recorded in several places, including [4, p. 71], [5, p. 83], [6, E31], and as Problem #423 on Bloom’s Erdős Problems website [2].
@@
In this paper we prove Conjecture 1.2. In fact, our argument applies to a much broader class of related greedy sequences obtained from an arbitrary finite initial seed.
Definition 1.3.
Let $s\geq 2$ and let $\mathbf{u}=(u_{1},\dots,u_{s})$ be a seed of length $s$ consisting of positive integers (not necessarily distinct or ordered). Define a sequence of positive integers $(a_{n}^{\mathbf{u}})_{n\geq 1}$ by
$a_{n}^{\mathbf{u}}:=u_{n}\qquad(1\leq n\leq s),$
and for $n\geq s+1$ let $a_{n}^{\mathbf{u}}$ be the least integer greater than $a_{n-1}^{\mathbf{u}}$ that can be written as a sum of at least two consecutive earlier terms:
$a_{n}^{\mathbf{u}}=\sum_{k=p}^{q}a_{k}^{\mathbf{u}}\qquad\text{for some }1\leq p\leq q\leq n-1,\ \ q-p\geq 1.$
We also set
$b_{n}^{\mathbf{u}}:=a_{n}^{\mathbf{u}}-n\qquad(n\geq 1).$
This definition is well defined: for each $n\geq s+1$, the sum $a_{n-2}^{\mathbf{u}}+a_{n-1}^{\mathbf{u}}$ is a valid candidate for $a_{n}^{\mathbf{u}}$, and it is strictly greater than $a_{n-1}^{\mathbf{u}}$ because all terms are positive. Therefore the minimum in the definition of $a_{n}^{\mathbf{u}}$ always exists. In particular, we have $a_{n+1}^{\mathbf{u}}>a_{n}^{\mathbf{u}}$ for all $n\geq s$.
When $\mathbf{u}=(1,2)$, the sequence $(a_{n}^{\mathbf{u}})_{n\geq 1}$ coincides with the classical sequence (1.1).
@@
For every integer $s\geq 2$ and every seed $\mathbf{u}$ of length $s$ consisting of positive integers, the sequence $(b_{n}^{\mathbf{u}})_{n\geq 1}$ is eventually nondecreasing; indeed, it is nondecreasing for all $n\geq s$. Moreover, $(b_{n}^{\mathbf{u}})_{n\geq 1}$ is unbounded.
The quantity $b_{n}^{\mathbf{u}}=a_{n}^{\mathbf{u}}-n$ measures the excess of the current term over its index. For all sufficiently large $n$, it gives a lower bound for the number of omitted positive integers in $1,a_{n}^{\mathbf{u}}]$. As a consequence, we obtain the following corollary.
Corollary 1.5.
For every integer $s\geq 2$ and every seed $\mathbf{u}$ of length $s$ consisting of positive integers,
$a_{n}^{\mathbf{u}}=n+\omega(1).$
In particular, the set $\{a_{n}^{\mathbf{u}}:n\geq 1\}$ omits infinitely many positive integers.
In particular, taking $\mathbf{u}=(1,2)$ in Corollary 1.5 proves Conjecture 1.2.
For the classical sequence (1.1), we also obtain a quantitative upper bound. In fact, the same argument extends to every fixed positive seed; see Remark 4.1.
@@
The proof of Theorem 1.6 combines the greedy structure of the sequence with a recent lower bound for difference sets of finite convex sets. Thus, besides proving that the sequence omits infinitely many positive integers, we also obtain a first polynomial upper bound toward Problem 1.1.
@@
We use Vinogradov’s asymptotic notation. For functions $f=f(n)$ and $g=g(n)$, we write $f=O(g)$, $f\ll g$, or $g\gg f$ to mean that there exists a constant $C>0$ such that $|f(n)|\leq Cg(n)$ for all sufficiently large $n$. If the implicit constant is allowed to depend on one or more parameters $z_{1},\dots,z_{r}$, we indicate this by writing $f\ll_{z_{1},\dots,z_{r}}g$, $g\gg_{z_{1},\dots,z_{r}}f$ or $f=O_{z_{1},\dots,z_{r}}(g)$. We write $f\asymp g$ or $f=\Theta(g)$ to mean that $f\ll g$ and $g\ll f$, and write $f=o(g)$ to mean $f(n)/g(n)\to 0$ as $n\to\infty$. We write $f=\omega(g)$ to mean that $g=o(f)$, equivalently, $|f(n)|/g(n)\to\infty$ when $n\to\infty$, assuming $g(n)>0$ for all sufficiently large $n$. In particular, $\omega(1)$ denotes any quantity tending to infinity as $n\to\infty$.
@@
Section 2 collects the preliminary ingredients, including the characterization of integers that can be written as sums of at least two consecutive positive integers and a Diophantine finiteness consequence of a result of Schinzel and Tijdeman. In Section 3 we prove Theorem 1.4 and Corollary 1.5. Section 4 is devoted to the proof of the polynomial upper bound in Theorem 1.6. Finally, Section 5 contains concluding remarks and numerical observations.
@@ 2. Preliminaries
Next we record the Diophantine ingredient. We shall use the following consequence of a result of Schinzel and Tijdeman [7, Corollary 1].
@@
has only finitely many integer solutions $(v,k)\in\mathbb{Z}\times\mathbb{Z}$ with $k>2$.
@@ 3. A qualitative lower bound and infinitely many omissions
Throughout this section, we fix a seed $\mathbf{u}$ of length $s\geq 2$, and write
$a_{n}:=a_{n}^{\mathbf{u}},\qquad b_{n}:=b_{n}^{\mathbf{u}}\qquad(n\geq 1).$
Our goal is to show that $(b_{n})$ is unbounded. The first step is a simple monotonicity observation, which will later force eventual linearity if $(b_{n})$ were bounded.
@@ 3.1. Monotonicity and eventual linearity
For every $n\geq s$, the greedy construction gives $a_{n+1}>a_{n}$, hence $a_{n+1}\geq a_{n}+1$. Therefore, for $n\geq s$,
@@
If $(b_{n})_{n\geq s}$ is bounded above, then there exist integers $n_{0}\geq s$ and $B$ such that
$a_{n}=n+B\qquad\text{for all }n\geq n_{0}.$
@@
By Lemma 3.1, the sequence $(b_{n})_{n\geq s}$ is a nondecreasing sequence of integers. If $(b_{n})_{n\geq s}$ is bounded above, then $(b_{n})_{n\geq s}$ must be eventually constant: there exist $n_{0}\geq s$ and an integer $B$ such that $b_{n}=B$ for all $n\geq n_{0}$. Hence $a_{n}=n+b_{n}=n+B$ for all $n\geq n_{0}$. ∎
From now on we assume, for contradiction, that $(b_{n})_{n\geq s}$ is bounded above and hence that
@@ 3.2. Consecutive-block decompositions beyond the linear tail
(Here $p=n_{0}$ contributes the empty sum $0\in\mathcal{C}$.) The next lemma isolates the only two ways a large term can be written as a sum of consecutive earlier terms once the tail is linear.
@@ 3.3. Powers of two force a quadratic exponential equation
Fix $r$ satisfying the inequality (3.2) and set $n=2^{r}-B$. Then $n\geq n_{0}$ and (3.1) gives $a_{n}=n+B=2^{r}$. Moreover, $n\geq s+1$, so by construction of the sequence there exist indices $1\leq p\leq q\leq n-1$ with $q-p\geq 1$ and
@@ 3.4. Completion of the proof
Lemma 3.1 shows that $(b_{n})$ is nondecreasing for all $n\geq s$. It remains to prove that it is unbounded.
Assume, for contradiction, that $(b_{n})$ is bounded above. Then (3.1) holds by Lemma 3.2. Lemma 3.4 shows that for every sufficiently large $r$ we can find $C\in\mathcal{C}$ and $v\geq T$ such that (3.3) holds. The set
@@
is finite. Hence by the pigeonhole principle there exists some $E\in\mathcal{E}$ for which the equation
@@
has infinitely many integer solutions $(v,k)\in\mathbb{Z}\times\mathbb{Z}$ with $k>2$. This contradicts Corollary 2.3. Therefore $(b_{n})$ is unbounded, as claimed. ∎
Proof of Corollary 1.5.
By Theorem 1.4, the sequence $(b_{n})$ is nondecreasing for all $n\geq s$ and unbounded, hence $b_{n}\to\infty$. Therefore
@@
Let $M:=\max_{1\leq i\leq s}a_{i}$. For all sufficiently large $n$, we have $n\geq s$ and $a_{n}\geq M$. Hence every value attained by the sequence in $1,a_{n}]$ occurs among the first $n$ terms: indeed, the initial seed values are all $\leq M\leq a_{n}$, while $a_{m}>a_{n}$ for every $m>n$ because the sequence is strictly increasing from index $s$ onward. The set of distinct values among $a_{1},\dots,a_{n}$ therefore has cardinality at most $n$, so the number of positive integers in $1,a_{n}]$ omitted by the sequence is at least
$a_{n}-n=b_{n}.$
@@
Since $b_{n}\to\infty$, the number of omitted integers up to $a_{n}$ is unbounded. In particular, the set $\{a_{n}:n\geq 1\}$ omits infinitely many positive integers. ∎
4. A quantitative upper bound for the classical sequence
In this section we restrict attention to the classical sequence (1.1) and prove the bound
$a_{n}\ll_{\varepsilon}n^{4175/2506+\varepsilon}.$
Remark 4.1.
For every fixed seed $\mathbf{u}$ of length $s\geq 2$ consisting of positive integers, the same exponent can also be obtained for $a_{n}^{\mathbf{u}}$, namely $a_{n}^{\mathbf{u}}\ll_{\mathbf{u},\varepsilon}n^{4175/2506+\varepsilon}$. Indeed, by Definition 1.3, the sequence $(a_{n}^{\mathbf{u}})_{n\geq s}$ is strictly increasing, so one may apply the argument of the present section to the tail
$a_{s}^{\mathbf{u}},a_{s+1}^{\mathbf{u}},a_{s+2}^{\mathbf{u}},\dots$
after making the obvious notational changes in the definitions of the relevant prefix sums and representable integers. Since this introduces additional bookkeeping but no essentially new idea, we present the details only for the classical seed $(1,2)$.

### 23 Mar 2026 v2

@@ 1. Introduction
A self-generating greedy process, often discussed under the name Hofstadter sequence (see, for instance, [14] and [8, E31]), starts from a short initial list and repeatedly adjoins sums of consecutive earlier terms. In the classical case one begins with $(1,2)$ and defines a strictly increasing sequence $(a_{n})_{n\geq 1}$ by
$a_{1}=1,\qquad a_{2}=2,$
and, for $k\geq 3$, taking $a_{k}$ to be the least integer greater than $a_{k-1}$ that can be written as a sum of at least two consecutive earlier terms:
@@
This is sequence A005243 in the OEIS [11]. The asymptotic behavior of $(a_{n})_{n\geq 1}$ was asked about by Hofstadter and has been recorded in several places, including [6, p. 71], [7, p. 83], [8, E31], and as Problem #423 on Bloom’s Erdős Problems website [5].
@@
For later use, write
$b_{n}:=a_{n}-n\qquad(n\geq 1).$
Our first result settles Conjecture 1.2.
@@
The sequence $(b_{n})_{n\geq 1}$ is nondecreasing and unbounded. Consequently,
$a_{n}=n+\omega(1),$
and the set $\{a_{n}:n\geq 1\}$ omits infinitely many positive integers.
In fact, building on Theorem 1.3, we prove the following quantitative lower bound.
Theorem 1.4.
We have
$a_{n}\geq n+\frac{\log\log n}{\log 20}-O(1).$
We also obtain a quantitative upper bound.
@@
The proof of Theorem 1.5 combines the greedy structure of the sequence with a recent lower bound for difference sets of finite convex sets. Thus, besides settling Conjecture 1.2, we also obtain a first polynomial upper bound toward Problem 1.1. For related work on consecutive sums in integer sequences, see Beker [1], Hegyvári [9], and Tao [13].
@@
We use Vinogradov’s asymptotic notation. For functions $f=f(n)$ and $g=g(n)$, we write $f=O(g)$, $g=\Omega(f)$, $f\ll g$, or $g\gg f$ to mean that there exists a constant $C>0$ such that $|f(n)|\leq Cg(n)$ for all sufficiently large $n$. If the implicit constant is allowed to depend on one or more parameters $z_{1},\dots,z_{r}$, we indicate this by writing $f=O_{z_{1},\dots,z_{r}}(g)$, $g=\Omega_{z_{1},\dots,z_{r}}(f)$, $f\ll_{z_{1},\dots,z_{r}}g$ or $g\gg_{z_{1},\dots,z_{r}}f$. We write $f\asymp g$ or $f=\Theta(g)$ to mean that $f\ll g$ and $g\ll f$, and write $f=o(g)$ to mean that $f(n)/g(n)\to 0$ as $n\to\infty$. We write $f=\omega(g)$ to mean that $g=o(f)$; in particular, $\omega(1)$ denotes any quantity tending to infinity as $n\to\infty$.
@@
Section 2 collects the two preliminary ingredients used later: the characterization of sums of consecutive positive integers and a Diophantine finiteness consequence of the theorem of Schinzel and Tijdeman. Section 3 proves Theorem 1.3. Section 4 establishes the quantitative lower bound in Theorem 1.4. Section 5 proves the polynomial upper bound in Theorem 1.5. Finally, Section 6 contains concluding remarks on extensions, later improvements, and open questions.
@@ 2. Preliminaries
Next we record the Diophantine ingredient. We shall use the following consequence of a result of Schinzel and Tijdeman [10, Corollary 1].
@@
has only finitely many integer solutions $(v,k)\in\mathbb{Z}^{2}$ with $k>2$.
@@ 3. A qualitative lower bound and infinitely many omissions
Throughout this section we write
$b_{n}:=a_{n}-n\qquad(n\geq 1).$
Our goal is to prove Theorem 1.3. The first step is a simple monotonicity observation.
@@ 3.1. Monotonicity and eventual linearity
For every $n\geq 1$, the greedy construction gives $a_{n+1}>a_{n}$, hence $a_{n+1}\geq a_{n}+1$. Therefore
@@
If $(b_{n})_{n\geq 1}$ is bounded above, then there exist integers $n_{0}\geq 2$ and $B$ such that
$a_{n}=n+B\qquad(n\geq n_{0}).$
@@
By Lemma 3.1, the sequence $(b_{n})_{n\geq 1}$ is a nondecreasing sequence of integers. If it is bounded above, then it is eventually constant: there exist $n_{0}\geq 2$ and an integer $B$ such that $b_{n}=B$ for all $n\geq n_{0}$. Hence $a_{n}=n+b_{n}=n+B$ for all $n\geq n_{0}$. ∎
From now on we assume, for contradiction, that $(b_{n})_{n\geq 1}$ is bounded above and hence that
@@ 3.2. Consecutive-block decompositions beyond the linear tail
(Here $p=n_{0}$ contributes the empty sum $0\in\mathcal{C}$.) The next lemma isolates the only two possible shapes of a large consecutive-block representation once the tail is linear.
@@ 3.3. Powers of two force a quadratic exponential equation
Fix $r$ satisfying the inequality (3.2) and set $n=2^{r}-B$. Then $n\geq n_{0}$ and (3.1) gives $a_{n}=n+B=2^{r}$. Moreover, $n\geq 3$, so by construction of the sequence there exist indices $1\leq p\leq q\leq n-1$ with $q-p\geq 1$ and
@@ 3.4. Completion of the proof
Lemma 3.1 shows that $(b_{n})$ is nondecreasing. It remains to prove that it is unbounded.
Assume, for contradiction, that $(b_{n})$ is bounded above. Then (3.1) holds by Lemma 3.2. By Lemma 3.4, for every sufficiently large $r$ there exist $C\in\mathcal{C}$ and $v\geq T$ such that (3.3) holds. The set
@@
is finite. Hence, by the pigeonhole principle, there exists some $E\in\mathcal{E}$ for which the equation
@@
has infinitely many integer solutions $(v,k)\in\mathbb{Z}^{2}$ with $k>2$. This contradicts Corollary 2.3. Therefore $(b_{n})$ is unbounded.
Since $(b_{n})$ is a nondecreasing unbounded sequence of integers, we have $b_{n}\to\infty$, and hence
@@
Finally, because the sequence $(a_{n})$ is strictly increasing, exactly $n$ elements of $\{a_{m}:m\geq 1\}$ lie in the interval $1,a_{n}]$, namely $a_{1},\dots,a_{n}$. Therefore at least
$a_{n}-n=b_{n}$
@@
positive integers in $1,a_{n}]$ are omitted by the sequence. As $b_{n}\to\infty$, the set $\{a_{n}:n\geq 1\}$ omits infinitely many positive integers. ∎
4. A quantitative lower bound for the classical sequence
The following explicit consequence of Beukers’ work [2, Corollary 1] will be used.
Lemma 4.1 ([2]).
Let $D\in\mathbb{Z}\setminus\{0\}$. Every integer solution $(x,m)\in\mathbb{Z}^{2}$ of
$x^{2}+D=2^{m}$
satisfies
$m<435+\frac{10\log|D|}{\log 2}.$
Lemma 4.2.
Let $\widehat{B}\geq 1$, and suppose that
$b_{n}=\widehat{B}\qquad(n_{1}\leq n\leq n_{2})$
for some integers $2\leq n_{1}\leq n_{2}$. Set
$\widehat{T}:=n_{1}+\widehat{B},\qquad\widehat{S}:=\sum_{i=1}^{n_{1}-1}a_{i}.$
If $r\in\mathbb{Z}$ satisfies
$\max\{\widehat{T},\ \widehat{S}+1,\ \widehat{B}+3\}\leq 2^{r}\leq n_{2}+\widehat{B}$, (4.1)
then there exist integers $x$ and $D\neq 0$ such that
$2^{r+3}=x^{2}+D\qquad\text{and}\qquad|D|\leq 13\widehat{T}^{2}.$
Proof.
Since $(b_{n})_{n\geq 1}$ is nondecreasing by Lemma 3.1 and $b_{n_{1}}=\widehat{B}$, we have
$b_{i}\leq\widehat{B}\qquad(1\leq i<n_{1}).$
Hence
$\widehat{S}=\sum_{i=1}^{n_{1}-1}a_{i}=\sum_{i=1}^{n_{1}-1}(i+b_{i})\leq\sum_{i=1}^{n_{1}-1}(i+\widehat{B})\leq\widehat{T}^{2}.$
Now fix $r$ satisfying (4.1), and set $\widehat{n}:=2^{r}-\widehat{B}$. Then
$n_{1}\leq\widehat{n}\leq n_{2},$
so $a_{\widehat{n}}=\widehat{n}+\widehat{B}=2^{r}$. Also $\widehat{n}\geq 3$, and therefore the greedy definition of the classical sequence (1.1) yields indices $1\leq p<q\leq\widehat{n}-1$ such that
$2^{r}=a_{\widehat{n}}=\sum_{k=p}^{q}a_{k}.$
Since $2^{r}>\widehat{S}$, we must have $q\geq n_{1}$. If $p\geq n_{1}$, then
$2^{r}=\sum_{k=p}^{q}(k+\widehat{B})=\sum_{j=p+\widehat{B}}^{q+\widehat{B}}j$
is a sum of at least two consecutive positive integers, contradicting Lemma 2.1. Hence $p<n_{1}\leq q$.
Set
$\widehat{C}:=\sum_{k=p}^{n_{1}-1}a_{k},\qquad\widehat{v}:=q+\widehat{B}$.
Then $0\leq\widehat{C}\leq\widehat{S}$ and
$2^{r}=\widehat{C}+\sum_{j=\widehat{T}}^{\widehat{v}}j=\widehat{C}+\frac{\widehat{v}(\widehat{v}+1)-\widehat{T}(\widehat{T}-1)}{2}.$
Thus
$2^{r+1}=\widehat{v}^{2}+\widehat{v}+2\widehat{C}-\widehat{T}(\widehat{T}-1),$
and therefore
$2^{r+3}=(2\widehat{v}+1)^{2}+D,\qquad D:=8\widehat{C}-4\widehat{T}(\widehat{T}-1)-1.$
Since $D\equiv-1\pmod{4}$, we have $D\neq 0$. Finally, using $0\leq\widehat{C}\leq\widehat{S}\leq\widehat{T}^{2}$, we obtain
$|D|\leq 8\widehat{C}+4\widehat{T}(\widehat{T}-1)+1\leq 8\widehat{T}^{2}+4\widehat{T}^{2}+1\leq 13\widehat{T}^{2}.$
This proves the lemma. ∎
We are now ready to prove the quantitative lower bound for the sequence (1.1).
Proof of Theorem 1.4.
By Theorem 1.3, the sequence $(b_{n})$ is nondecreasing and unbounded. Hence, for each integer $\widehat{B}\geq 0$,
$N(\widehat{B}):=\max\{n\geq 1:\ b_{n}\leq\widehat{B}\}$
is well defined and finite. Set
$M(\widehat{B}):=N(\widehat{B})+\widehat{B}+2\qquad(\widehat{B}\geq 0).$
We claim that there exists an absolute constant $K\geq 2$ such that
$M(\widehat{B})\leq K\,M(\widehat{B}-1)^{20}\qquad(\widehat{B}\geq 1).$ (4.2)
Fix $\widehat{B}\geq 1$. If $N(\widehat{B})=N(\widehat{B}-1)$, then
$M(\widehat{B})=N(\widehat{B})+\widehat{B}+2=N(\widehat{B}-1)+\widehat{B}+2=M(\widehat{B}-1)+1.$
Since $M(\widehat{B}-1)\geq 2$, it follows that
$M(\widehat{B})\leq M(\widehat{B}-1)^{20}.$
Suppose now that $N(\widehat{B})>N(\widehat{B}-1)$. Since $(b_{n})$ is integer-valued and nondecreasing, we then have
[…]
Write
$n_{1}:=N(\widehat{B}-1)+1,\qquad n_{2}:=N(\widehat{B}),\qquad\widehat{T}:=n_{1}+\widehat{B}$.
Then
$\widehat{T}=N(\widehat{B}-1)+1+\widehat{B}=N(\widehat{B}-1)+(\widehat{B}-1)+2=M(\widehat{B}-1).$
Let
$K_{0}:=2^{432}13^{10}.$
We shall show that
$n_{2}+\widehat{B}\leq 2K_{0}\widehat{T}^{20}.$
Assume to the contrary that
$n_{2}+\widehat{B}>2K_{0}\widehat{T}^{20}.$
Then there exists an integer $r$ such that
$K_{0}\widehat{T}^{20}<2^{r}\leq 2K_{0}\widehat{T}^{20}<n_{2}+\widehat{B}$.
We now verify the hypotheses of Lemma 4.2. Since $\widehat{T}\geq 2$, we have
$2^{r}>K_{0}\widehat{T}^{20}\geq\widehat{T}^{20}>\widehat{T}^{2}.$
Moreover, for every $1\leq i\leq n_{1}-1$ we have $b_{i}\leq\widehat{B}-1$, because $n_{1}=N(\widehat{B}-1)+1$. Hence
$a_{i}=i+b_{i}\leq i+\widehat{B}-1\qquad(1\leq i\leq n_{1}-1),$
and therefore
$\sum_{i=1}^{n_{1}-1}a_{i}\leq\sum_{i=1}^{n_{1}-1}(i+\widehat{B}-1)=\frac{(n_{1}-1)(n_{1}+2\widehat{B}-2)}{2}.$
On the other hand,
$\widehat{T}^{2}=(n_{1}+\widehat{B})^{2},$
and a direct calculation shows that
$2\widehat{T}^{2}-(n_{1}-1)(n_{1}+2\widehat{B}-2)=n_{1}^{2}+2n_{1}\widehat{B}+2\widehat{B}^{2}+3n_{1}+2\widehat{B}-2>0.$
Thus
$\sum_{i=1}^{n_{1}-1}a_{i}\leq\widehat{T}^{2}<2^{r}.$
Also, from $2^{r}>\widehat{T}^{2}$ and $\widehat{T}\geq 2$, we obtain
$2^{r}\geq\widehat{T}$.
Finally, since $n_{1}\geq 2$, we have
$\widehat{T}=n_{1}+\widehat{B}\geq\widehat{B}+2,$
and therefore
$2^{r}>\widehat{T}^{2}\geq(\widehat{B}+2)^{2}>\widehat{B}+3,$
so in particular
$2^{r}\geq\widehat{B}+3.$
Hence Lemma 4.2 applies, and yields integers $x$ and $D\neq 0$ such that
$2^{r+3}=x^{2}+D\qquad\text{and}\qquad|D|\leq 13\widehat{T}^{2}.$
Applying Lemma 4.1, we obtain
$r+3<435+\frac{10\log|D|}{\log 2}\leq 435+\frac{10\log(13\widehat{T}^{2})}{\log 2}.$
Exponentiating, we get
$2^{r}<2^{432}13^{10}\widehat{T}^{20}=K_{0}\widehat{T}^{20},$
contrary to the choice of $r$. This contradiction proves that
$n_{2}+\widehat{B}\leq 2K_{0}\widehat{T}^{20}.$
[…]
$N(\widehat{B})+\widehat{B}\leq 2K_{0}\,M(\widehat{B}-1)^{20},$
and hence
$M(\widehat{B})=N(\widehat{B})+\widehat{B}+2\leq(2K_{0}+2)\,M(\widehat{B}-1)^{20}.$
Thus (4.2) holds with
$K:=2K_{0}+2.$
Taking logarithms in (4.2) and iterating, we obtain
$\log M(\widehat{B})\leq 20^{\widehat{B}}\log M(0)+\bigl(1+20+\cdots+20^{\widehat{B}-1}\bigr)\log K$
for all $\widehat{B}\geq 0$, whence
$\log M(\widehat{B})\leq 20^{\widehat{B}}\log M(0)+\frac{20^{\widehat{B}}-1}{19}\log K\leq\widehat{A}\,20^{\widehat{B}}$,
where
$\widehat{A}:=\log M(0)+\frac{\log K}{19}.$
[…]
$\log n\leq\log M(\widehat{B})\leq\widehat{A}\,20^{\widehat{B}}$.
Taking logarithms once more, we get
$\widehat{B}=b_{n}\geq\frac{\log\log n-\log\widehat{A}}{\log 20}.$
Finally, since $a_{n}=n+b_{n}$, it follows that
$a_{n}\geq n+\frac{\log\log n}{\log 20}-\frac{\log\widehat{A}}{\log 20}.$
This proves the theorem. ∎
5. A polynomial upper bound for the classical sequence
In this section we prove Theorem 1.5. The argument combines the greedy ordering of the representable integers with a lower bound for difference sets of finite convex sets.